OpenStudy (anonymous):

Please help me! Will give medal :) The volume of a cone is 300 cubic kilometers. If the radius of the base is 7 kilometers, what is the height of the cone? Use three point one four for pi. Show your work and round to the nearest tenth.

OpenStudy (anonymous):

Please help! last question for my final exam... :)

OpenStudy (anonymous):

@nincompoop

OpenStudy (nincompoop):

do you know the formula of the volume of a cone?

OpenStudy (nincompoop):

\[V=\pi r^2 \frac{ h }{ 3 }\]

OpenStudy (anonymous):

i was just gonna write that lol

OpenStudy (nincompoop):

so solve for h \[\frac{ 3V }{ \pi r^2 }=h\]

OpenStudy (anonymous):

V = πr^2h 300 km^3 = 3.14(7 km)^2 h 1.949824516... = h The height is approximately 1.9 km

OpenStudy (nincompoop):

\[\frac{ 3\times 300 }{ \pi \times 7^2 }\]

OpenStudy (nincompoop):

neroye, please do not give out answers. it is a violation of the code of conduct

OpenStudy (anonymous):

3*300=900... 3.14*49=153.86

OpenStudy (nincompoop):

did you use a calculator? it says nearest tenth

OpenStudy (anonymous):

yes i used calculator.. my teacher says i can.. :)

OpenStudy (anonymous):

so 154

OpenStudy (anonymous):

153 or 154.. i think its 153

OpenStudy (nincompoop):

nope

OpenStudy (nincompoop):

did you look at the place values?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i reccommed not use a calculator it will ruin your math just an advice

OpenStudy (anonymous):

gotya chief :)

OpenStudy (nincompoop):

12345.6789 nearest tenth would be looking at the 6 and you have to look at the number to its right. if it is 5 or more you add 1 to 6

OpenStudy (anonymous):

you know how to round off right?

OpenStudy (nincompoop):

so that gives you 12345.7 as nearest tenth round off

OpenStudy (nincompoop):

@ChiefArnav I got this already. you can help other people

OpenStudy (anonymous):

i'am justing a job from my part know u didi it but i hiope u got me

OpenStudy (anonymous):

150.9

OpenStudy (anonymous):

do u want a re expalination

OpenStudy (anonymous):

umm.. its ok @nincompoop can help me :) thanks for your help though..

OpenStudy (anonymous):

okay

OpenStudy (nincompoop):

correct

OpenStudy (anonymous):

ok... so is that all?

OpenStudy (nincompoop):

yeah

OpenStudy (anonymous):

may be guess so

OpenStudy (nincompoop):

we've done a formula transformation that is we first determined the formula of h in a volume of a cone then use that formula to solve for the actual value with the information we were provided.

OpenStudy (anonymous):

got it! thank you all so much!

OpenStudy (nincompoop):

you're welcome